The Monty Company issued $240,000 of 13% bonds on January 1, 2017. The bonds are due January 1, 2022, with interest payable each July 1 and January 1. The bonds were issued at 96.

Prepare the journal entries for (a) January 1, (b) July 1, and (c) December 31. Assume The Monty Company records straight-line amortization semiannually. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Round intermediate calculations to 6 decimal places, e.g. 1.251247 and final answer to 0 decimal places, e.g. 38,548.)

No.

Date

Account Titles and Explanation

Debit

Credit

(a)

January 1, 2017

Entry field with correct answer

Entry field with incorrect answer now contains modified data

Entry field with correct answer

Entry field with correct answer

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with incorrect answer

(b)

Entry field with correct answer Jan. 1, 2017July 1, 2017Dec. 31, 2017

Entry field with correct answer

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer

Entry field with incorrect answer

(c)

Entry field with correct answer Jan. 1, 2017July 1, 2017Dec. 31, 2017

Entry field with correct answer

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer

Entry field with incorrect answer
Answer #1
Account Titles and Explanation Debit Credit
a January 1, 2017 Cash 230400 =240000*0.96
Discount on Bonds payable 9600
       Bonds payable 240000
b July 1, 2017 Interest expense 16560
       Discount on Bonds payable 960 =9600/10
       Cash 15600 =240000*13%*6/12
c Dec. 31, 2017 Interest expense 16560
       Discount on Bonds payable 960
       Interest payable 15600